Solve by substitution method. Thank you.
7m+n=49
m-3n=29

Answer by JBarnum(2044)

7m+n=49
m-3n=29 or m=29+3n
7(29+3n)+n=49
203+21n+n=49
203+22n=49
22n=-154
n=-7
m-3(-7)=29
m+21=29
m=8